Cashis11:

Ernie bought 3 12-packs Of regular cola, 4 12- packs of diet cola, and 2 6-packs of orange soda. How many cans did he buy?

DanGonz:

hmmmm

DanGonz:

he bought 96 can of soda you have to round

XxAngelGirlxX:

108 cans 3*12=36 4*12=48 2*6=12 3+48+12=108

Cashis11:

Thanks

XxAngelGirlxX:

please close this question when done <3 welcome